ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: Kutz, Nadja, Hoffmann, Tim, Kellendonk, Johannes, Technische Universität Berlin, Reshetikhin, Nicolai, UC Berkeley Discrete Nonlinear Toda Evolution We investigate the Poisson geometry for discrete nonlinear Toda systems. Their discrete evolution is given by a factorization relation in a semisimple Poisson Lie group. In particular a careful description of the corresponding symplectic leaves will be given.